OpenStudy (anonymous):

using implicit differentiation, find dy/dx given cos(xy)=y^2. implicit differentiation is not exactly my strong spot :(

OpenStudy (dumbcow):

just remember when you take derivative of "y" term to multiply by "dy/dx" for y^2 , derivative is 2y times dy/dx for cos(xy) use chain rule and product rule

OpenStudy (anonymous):

thank you!!

OpenStudy (dumbcow):

yw you should get: \[-\sin(xy) (y + x \frac{dy}{dx}) = 2y \frac{dy}{dx}\] use algebra to isolate dy/dx